### v1

# Code Generation with Small Language Models: A Deep Evaluation on Codeforces

Abstract
Large Language Models (LLMs) have demonstrated capabilities in code generation, potentially boosting developer productivity. However, their widespread adoption remains limited by high computational costs, significant energy demands, and security risks such as data leakage and adversarial attacks. As a lighter-weight alternative, Small Language Models (SLMs) offer faster inference, lower deployment overhead, and better adaptability to domain-specific tasks, making them an attractive option for real-world applications. While prior research has benchmarked LLMs on competitive programming tasks, such evaluations often focus narrowly on metrics like Elo scores or pass rates, overlooking deeper insights into model behavior, failure patterns, and problem diversity. Furthermore, the potential of SLMs to tackle complex tasks such as competitive programming remains underexplored. In this study, we benchmark five open SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—across 280 Codeforces problems spanning Elo ratings from 800 to 2100 and covering 36 distinct topics. All models were tasked with generating Python solutions. Phi-4 14B achieved the best performance among SLMs, with a pass@3 of 63.6%, approaching the proprietary o3-mini-high (86.8%). In addition, we evaluated Phi-4 14B on C++ and found that combining outputs from both Python and C++ increases its aggregated pass@3 to 73.6%. A qualitative analysis of Phi-4 14B’s incorrect outputs revealed that some failures were due to minor implementation issues—such as handling edge cases or correcting variable initialization—rather than deeper reasoning flaws.
Index Terms:
Small Language Models, Code Generation, Competitive Programming.
I Introduction
Automatic code generation has become a strategic solution to meet rising demands for developer productivity and address the ongoing shortage of skilled professionals. Recent advances in this area have progressed from early approaches like natural language-based synthesis [1] and grammar-based models [2] to large pre-trained transformers such as CodeBERT [3] and Codex [4], which have significantly advanced the state of the art.
Despite these successes, benchmarks like HumanEval [4] primarily target short, simple problems [5, 6], limiting their applicability to real-world scenarios. More recent efforts have adopted competitive programming platforms such as Codeforces [7, 8], but often emphasize high-level metrics (e.g., Elo score, pass rate) without deeper analysis of problem diversity, model behavior, or failure modes. Additionally, evaluations based on offline datasets—such as those used by AlphaCode [9, 10]—may overlook key constraints like runtime and memory usage.
While Large Language Models (LLMs) achieve strong performance [6, 11, 12], their high computational demands, energy usage, and security risks—such as data leakage and intellectual property exposure [13, 14, 15, 16]—can hinder adoption. In contrast, Small Language Models (SLMs) present a promising alternative: with fewer parameters, they offer lower latency, easier deployment, and greater adaptability for fine-tuning in constrained environments [17]. However, their capabilities in domains like competitive programming remain underexplored.
In this study, we evaluate five general-purpose SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—on 280 Codeforces problems ranging from Elo 800 to 2100 and covering 36 topics, including Dynamic Programming, Strings, Segment Trees. All models were tasked with generating Python code, with three submissions per problem, evaluated using the Codeforces judge for correctness, runtime, and memory usage. Among the SLMs, Phi-4 14B achieved the highest performance, with a pass@3 of 63.6%. It was particularly effective on problems rated 800–1500 and in topics such as Strings (75.9%), Sorting (76.3%), and Mathematics (68.6%). It also demonstrated high semantic consistency, solving 77.5% of problems correctly within three attempts. In contrast, other SLMs—DeepSeek-R1 14B, Llama 3.2 3B, Gemma 3 12B, and Gemma 2 9B—performed significantly lower, each scoring below 24.0%.
We also evaluated Phi-4 14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@3 to 73.6%. For comparison, the proprietary o3-mini-high model achieved a pass@3 of 86.8%. A manual analysis of Phi-4 14B’s incorrect outputs, supported by an expert in competitive programming, revealed that many failures were caused by minor implementation issues, or language-specific limitations, rather than fundamental reasoning flaws. These results reinforce the practical potential of small, open models as efficient and reliable alternatives for code generation, approaching the performance of state-of-the-art proprietary solutions. All data and code are publicly available online [18].
II Methodology
Our study evaluates the effectiveness of open SLMs in solving programming problems commonly found on competitive programming platforms. We evaluate five open SLMs: Llama 3.2 3B [19], Gemma 2 9B [20], Gemma 3 12B [21], Phi-4 14B [22], and DeepSeek-R1 14B [23]. These models were selected based on their potential in code-centric tasks [24, 25, 23]. Additionally, they represent a diverse set of architectural designs and development teams.
To evaluate model performance, we selected 280 problems from the Codeforces, a widely used platform in education, technical interviews, and algorithmic competitions, known for its diverse and challenging problem sets. Its broad range of topics and difficulty levels makes it a strong benchmark for assessing the reasoning and code-generation capabilities of language models [26, 9, 7]. Solving Codeforces problems requires models to comprehend complex natural language descriptions, apply suitable algorithms and data structures, and produce syntactically and semantically correct implementations. Submissions are rigorously evaluated using hidden test suites that assess correctness, edge case handling, execution time, and memory usage [10]. Although the focus is on Codeforces programming problems, many of these tasks reflect real-world scenarios encountered by software engineers, involving the interpretation of requirement documents, the use of appropriate data structures, and the application of algorithms commonly used in day-to-day practice.
Problem Watermelon Description
One hot summer day Pete and his friend Billy decided to buy a watermelon. […] Pete and Billy are great fans of even numbers, that’s why they want to divide the watermelon in such a way that each of the two parts weighs even number of kilos, at the same time it is not obligatory that the parts are equal. […] For sure, each of them should get a part of positive weight.
Input
The first (and the only) input line contains integer number $w$ $(1\leq w\leq 100)$ — the weight of the watermelon bought by the boys.
Output
Print YES, if the boys can divide the watermelon into two parts, each of them weighing even number of kilos; and NO in the opposite case.
Examples
Input 8 Output YES
Note
For example, the boys can divide the watermelon into two parts of 2 and 6 kilos respectively (another variant — two parts of 4 and 4 kilos).
Codeforces ranks problems using an ELO-based system, with ratings ranging from 800 to over 3000. We focused on problems rated between 800 and 2100, as fewer than 2% of active Codeforces users had ratings above 2100 at the time of writing. For each rating level within this range, we selected 20 problems based on submission popularity as of December 15, 2024. Problem statements ranged from 111 to 1,117 tokens, with an average length of 432 and a median of 404 tokens. The dataset spans 36 topics, including Implementation, Mathematics, Dynamic Programming, and Data Structures. Implementation was the most frequent tag, typically associated with problems that are conceptually simple but require intricate coding. As an example, next we show part of the “Watermelon” problem [^1] https://codeforces.com/problemset/problem/4/A illustrating the typical structure of a Codeforces task, which includes a statement, input/output format, examples, and an optional note.
Codeforces problem statements are often indirect, embedding the task within a narrative that can obscure the core requirements. This design encourages competitors to reflect, understand the problem, and identify the appropriate algorithm or data structure. Such formulations can pose additional challenges for foundation models, which must accurately interpret implicit requirements.
We developed an automated pipeline to evaluate all SLMs. Each model submitted three solutions per problem. All executions were performed locally in March 2025 using the Ollama framework [27] on a NVIDIA RTX 3060 GPU (12GB VRAM), with default parameters from the LangChain Ollama API [28] to ensure consistency across models. The process was fully automated: problem statements were extracted, prompts generated, code produced by the models, and solutions submitted to Codeforces.
To improve solution accuracy, we designed prompts with essential contextual information, including a persona definition, language-specific instructions (Python), and clear guidelines on input/output handling. Additionally, the prompt describes the structure of the Codeforces problem statement, ensuring the model interprets it correctly. Together with the actual problem, these components provide a well-rounded foundation to guide the code generation process.
{mdframed}
[ backgroundcolor=cyan!10, linewidth=1pt, linecolor=black, innerleftmargin=5pt, innerrightmargin=5pt, ] You are a highly skilled competitive programmer with 15 years of experience in the field.
Your objective is to analyze the following problem statement and to produce a Python code solution that adheres to the requirements.
Guidelines for the solution: deliver only the Python code; Ensure the solution reads input via standard input and produces outputs results via standard output; If the solution requires defining a function, ensure it is executed within the code; Avoid adding explanations, comments, or unnecessary text.
The Problem Statement includes a detailed description, input and output format, and examples to clarify requirements.
{Problem Statement}
III Evaluation
We evaluate models using the pass@k metric [4], which estimates the probability that at least one of the top-k generated solutions is correct. As shown in Table I, Phi-4 14B (Python) outperforms all other models, achieving 48.3% at pass@1 and 63.6% at pass@3—nearly three times higher than the next best model, DeepSeek-R1 14B (15.5% at pass@1, 23.9% at pass@3). Gemma 3 12B follows with 16.9% and 19.6%, respectively, outperforming both Gemma 2 9B and Llama 3.2 3B, which stay below 11% at pass@3, suggesting that smaller models struggle with more complex programming tasks. We also analyzed the length of generated outputs. As shown in Table I, Phi-4 14B and Gemma 3 12B produce longer solutions, averaging 24.3 and 27.7 lines of code, with over 200 tokens on average. In contrast, DeepSeek-R1 14B generates more concise outputs (9.2 LOC, 82.2 tokens). Notably, Phi-4 14B also produces the longest individual solution (1,163 tokens).
TABLE I: Evaluation of Llama 3.2 3B (Lla), Gemma 2 9B (Gem2), Gemma 3 12B (Gem3), DeepSeek-R1 14B (DS), and Phi-4 14B (Phi) across pass@k, Codeforces topics, and semantic consistency.
Lla Gem2 Gem3 DS Phi
Accuracy (pass@k)
pass@1 7.0% 8.4% 16.9% 15.5% 48.3%
pass@2 9.6% 9.8% 18.6% 21.1% 58.8%
pass@3 11.1% 10.4% 19.6% 23.9% 63.6%
Topical Accuracy (pass@3)
Implementation 17.6% 20.9% 34.1% 41.3% 83.5%
Sortings 7.9% 2.6% 10.5% 18.9% 76.3%
Strings 34.5% 37.9% 51.7% 48.3% 75.9%
Two Pointers 0.0% 4.3% 4.3% 12.5% 52.2%
Math 8.1% 7.0% 19.8% 20.5% 68.6%
Brute Force 14.0% 10.0% 32.0% 28.0% 68.0%
Combinatorics 0.0% 0.0% 11.1% 33.3% 66.7%
Greedy 10.7% 4.0% 12.0% 13.3% 65.3%
Binary Search 0.0% 2.6% 5.1% 12.2% 61.5%
Constructive Alg. 3.0% 6.1% 3.0% 6.1% 57.6%
Semantic Consistency
Score 61.3% 86.2% 85.4% 64.2% 77.5%
Generated Code Length
LOC_MEAN 19.5 13.4 27.7 9.2 24.3
LOC_MAX 81 40 99 88 106
MEAN_TOKEN 164.8 102.7 210.2 82.2 204.7
MAX_TOKEN 752 431 762 917 1,163
Phi-4 in C++ (pass@1 / 2 / 3)
Phi-4 14B (C++) 47.7% / 57.3% / 61.1%
Accuracy of Additional Models (pass@1 / 2 / 3)
o3-mini-high 83.3% / 85.7% / 86.8%
III-A Types of Errors
When submitting a solution to a Codeforces problem, verdicts other than Accepted—such as Wrong Answer, Compilation Error, Time Limit Exceeded (TLE), Memory Limit Exceeded (MLE), and Runtime Error—indicate specific types of failures. TLE and MLE indicate that the submitted code fails to meet the time and memory constraints, respectively. Figure 1 presents the distribution of these errors across models. Wrong Answer is the most frequent error across all models, indicating incorrect outputs for one or more test cases. Runtime Error, often caused by invalid memory access or uninitialized variables, is the second most common failure, particularly affecting Llama 3.2 3B (27%) and Gemma 2 9B (15%). TLE is also notable, especially for Gemma 3 12B (39.3%) and Gemma 2 9B (10%), suggesting inefficiencies in the generated algorithms.
DeepSeek-R1 14B also had the lowest Wrong Answer rate (24.9%). Compilation Errors are especially high in DeepSeek-R1 14B (16.4%), while MLE is rare across all models. DeepSeek-R1 14B exhibited an unusual behavior, with 35.5% of its submissions labeled as Not Answered—a custom category for outputs that could not be submitted as valid code, such as explanations or partial snippets. This indicates a difficulty in generating structured, submission-ready solutions, which limits its usability in automated pipelines.
Llama 3.2 3BGemma 2 9BGemma 3 12BPhi-4 14BDeepSeek-R1 14B0020202020404040406060606058.558.558.558.566.266.266.266.253.953.953.953.933.933.933.933.924.924.924.924.927272727151515153.63.63.63.66.46.46.46.47.67.67.67.655551010101039.339.339.339.39.49.49.49.40.40.40.40.42.52.52.52.50.50.50.50.51.41.41.41.40.20.20.20.216.416.416.416.40.20.20.20.2001.91.91.91.91.71.71.71.7000000000035.535.535.535.5Error Percentage (%)Wrong AnswerRuntime ErrorTime Limit ExceededCompilation ErrorMemory Limit ExceededNot Answered
Figure 1: Percentage of submissions that resulted in an error across 840 submissions (280 problems, each submitted three times).
III-B Model Effectiveness by Topic
Each Codeforces problem is associated with one or more topics. Table I summarizes model performance across 10 of the 36 topics considered in our study, including Math and Greedy algorithms—each requiring distinct forms of algorithmic reasoning. Phi-4 14B outperforms all other models across every topic, achieving the highest pass@3 accuracy. Its performance is particularly strong in Implementation (83.5%), Sortings (76.3%), and Strings (75.9%), with some gaps over the second-best model. It also leads in more complex categories like Binary Search (61.5%) and Brute Force (68%), demonstrating generalization across a range of problem types. DeepSeek-R1 14B ranks second overall. Gemma 3 12B consistently improves over Gemma 2 9B, especially in Strings (51.7%) and Brute Force (32%), suggesting that moderate increases in model size or training can yield measurable gains. Llama 3.2 3B and Gemma 2 9B struggle on more demanding topics. For instance, in Combinatorics, Llama 3.2 3B scores 0%, while Phi-4 14B achieves 66.7%. A similar trend appears in Binary Search, Constructive Algorithms, and Two Pointers, where only Phi-4 14B surpasses 50% accuracy.
III-C Performance Breakdown by Difficulty
As shown in Figure 2, Phi-4 14B (Python) consistently outperforms all open-small language models across nearly all difficulty levels. However, all SLMs show a performance decline as problem difficulty increases, highlighting their limitations with complex tasks and relative strength on easier ones. Each SLM exhibits different performance profiles. Llama 3.2 3B performs reasonably well up to rating 900, with only sporadic success beyond that. Gemma 2 9B shows moderate accuracy between 800–900, while Gemma 3 12B extends that range to around 1000, although with noticeable drops in accuracy after 1200. DeepSeek-R1 14B maintains relatively stable performance up to 1300. Phi-4 14B spans the full range up to 2100, but shows a decline beyond 1500, indicating challenges with higher complexity.
0.80.911.11.21.31.41.51.61.71.81.922.10%20%40%60%80%100%Codeforces Rating (K)Correct answers (%)Phi-4 14B (Python)DeepSeek-R1 14BGemma 3 12BLlama 3.2 3BGemma 2 9Bo3-mini-highPhi-4 14B (C++)Phi-4 (C++ and Python)
Figure 2: Performance of models on Codeforces problems across difficulty levels (800–2100). The y-axis indicates the percentage of correct solutions out of 20 problems per level.
III-D Consistency Analysis
While pass@k measures how often a model produces at least one correct solution, it does not capture consistency across multiple attempts. To address this, we use Semantic Consistency (SC), which evaluates how reliably a model reproduces correct outputs in repeated submissions [29]. This metric is particularly important for code generation, where consistent correctness matters more than isolated success. As shown in Table I, all models achieved SC above 60%, with Gemma 2 9B and Gemma 3 12B leading despite lower overall accuracy. Phi-4 14B also showed strong consistency (77.5%), indicating that once a model solves a problem, it tends to do so repeatedly—a sign of learned, stable behavior rather than random chance.
III-E Phi-4 14B: C++
To investigate whether the choice of programming language influences model performance, we evaluated Phi-4 14B, which has the best performance in our evaluation, using both Python and C++. C++ is widely adopted in competitive programming due to its performance efficiency, particularly in time-critical problems. In this experiment, we kept the prompting strategy, modifying only the target language.
The results show that Phi-4 14B achieves comparable accuracy in both languages: 63.6% pass@3 in Python and 61.1% in C++. However, the sets of problems solved in each language differ around 20%. The Python version solved 35 problems not solved in C++, while the C++ version succeeded on 28 problems not handled in Python. When combining both sets (see Figure 2), Phi-4 14B solves 206 out of 280 problems—achieving an aggregated pass@1 of 63.6%, pass@2 of 70.4%, and pass@3 of 73.6%.
Of the 28 problems uniquely solved in C++, 25 are rated above 1500 and primarily fall under categories such as Dynamic Programming, Math, Greedy, Brute Force, and Graphs. In these cases, Python submissions often failed due to Time Limit Exceeded (12), Wrong Answers (18), or Runtime Errors (9). These findings suggest that problem complexity and computational constraints may favor C++ over Python in certain scenarios.
[…]
III-F Larger Models
To understand how far SLMs are from the best-performing models, we analyzed OpenAI’s proprietary o3-mini-high, which, as shown in Table I, achieved the highest overall accuracy with a pass@3 of 86.8%. Phi-4 14B successfully solved five problems that were not answered by o3-mini-high. These problems had Elo ratings of 1400, 1500, 1900 (2x) and 2100. To further explore its limitations, we submitted the 38 problems that o3-mini-high failed to solve to OpenAI’s larger reasoning model, o1 [7], using three attempts per problem. These problems ranged from Elo 1400 to 2100 and focused on challenging categories such as Dynamic Programming, Math, Trees, and Data Structures. o1 successfully solved only 3 of the 38 problems (7.9% pass@3). Among the three successful cases, problem 479-C (Elo 1400) was solved by both o1 and Phi-4 14B, but not by o3-mini-high. The other two solutions involved complex Data Structures at Elo 2100. For the remaining 35 problems, the failures were mainly due to Time Limit Exceeded (30 cases), Wrong Answers (14), Memory Limit Exceeded (5), and one Runtime Error. These findings confirm that even state-of-the-art proprietary models struggle with high-difficulty competitive programming tasks.
However, when aggregating Phi-4 14B’s results across Python and C++ (Section III-E), its pass@3 increases to 73.6%, narrowing the gap to o3-mini-high. This highlights the potential of open models to achieve competitive performance through multi-language strategies. Both models follow a similar trend: strong accuracy on easier problems with a gradual decline as difficulty increases (Figure 2). Moreover, Phi-4 14B is open, cost-effective, and runs on consumer-grade hardware, making it an attractive alternative for a wide range of applications.
[…]
III-G Qualitative Analysis
To complement our quantitative evaluation metrics such as pass@k and semantic consistency, we conducted a qualitative analysis of the errors made by Phi-4 14B. One of the authors, a competitive programmer with over seven years of experience, has reached a peak ELO rating of 2245 on Codeforces. He conducted a manual review of a sample of 21 (~$\texttildelow 20\%$) problems that Phi-4 14B failed to solve. Based on this analysis, we identified four primary causes of failure: near-correct solutions with minor mistakes, algorithmically correct solutions that exceeded time limits, incorrect solutions with relevant direction, and fundamentally flawed implementations.
The first category of failures involved near-correct solutions that required only minor adjustments to succeed. In five of the reviewed problems, Phi-4 14B produced almost correct code, needing simple fixes such as correcting variable initialization (Problem 1343-C), handling edge cases (Problem 1399-C), or adjusting state transitions in dynamic programming (Problem 698-A). Three additional problems required more modifications. In Problem 489-C, multiple assignment errors and output formatting issues had to be corrected. Problem 1-B involved a minor logical issue, which required adding a loop to properly parse the input. In Problem 161-D, the Python solution was logically sound but inefficient; performance improvements were necessary, including replacing dictionaries with lists and rewriting recursive functions as iterative ones to meet time constraints.
In the second category, three problems featured conceptually correct algorithms that failed due to time constraints. When we prompted Phi-4 14B to rewrite these solutions in C++—preserving the logic but aiming for improved efficiency—two were successfully solved. However, Problem 1541-B (Elo 1200) remained unsolved, as it required a reduction in time complexity beyond what the model could infer.
[…]
In 4 of the reviewed problems, Phi-4 14B produced outputs that were in the right direction but still far from being correct. These solutions often included algorithmic elements aligned with the correct approach—such as identifying the right data structure or outlining the high-level logic—but lacked crucial implementation details or contained major logical errors. While the model appeared to grasp the core idea behind the problems, its execution was flawed and incomplete. As a result, these attempts were not easily repairable and required substantial rewriting to become viable. Still, one noteworthy aspect is that this type of partially correct attempt can offer valuable insights to the user. In problems where it is particularly hard to identify a viable direction, having the model suggest an initial high-level approach—even if incomplete—can be quite helpful as a starting point for further reasoning. In Problem 339-D (Elo 1700), the failure was attributed to incomplete problem description caused by formatting issues during automatic extraction. For example, numbers with exponents were incorrectly ignored, altering the meaning of the problem statement. As future work, we plan to enhance the text extraction process to avoid such issues.
Finally, in 5 of the 21 cases, the model produced fundamentally incorrect code. These outputs were either logically incoherent or based on an incorrect understanding of the problem statement. In some cases, the specialist reviewer noted that “the SLM seems not to have understood the problem at all.” In all three attempts, the model made incorrect assumptions and attempted to solve a completely different task.
[…]
IV Conclusion
This study evaluated the performance of five open SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—on 280 competitive programming problems from Codeforces. Phi-4 14B stood out with a pass@3 of 63.6% and semantic consistency of 77.5%, significantly outperforming other SLMs. In contrast, smaller models like Llama 3.2 3B and Gemma 2 9B achieved pass@3 rates below 24%. We also evaluated Phi-4 14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@3 to 73.6%. For comparison, the proprietary o3-mini-high model achieved a pass@3 of 86.8%.
A manual review of incorrect Phi-4 14B submissions revealed that most solutions were close to correct, with errors arising from minor issues—such as unhandled edge cases or suboptimal implementations—rather than fundamental reasoning flaws. Notably, 26 problems that failed in Python were successfully solved in C++, suggesting that leveraging multiple programming languages can enhance model performance. These results show that SLMs—particularly Phi-4 14B—strike a compelling balance between performance, efficiency, and accessibility. They represent a practical alternative to large proprietary models, especially in settings with limited resources.
While previous works have analyzed the performance of models like AlphaCode [9] and o3-mini-high [7] on Codeforces, focusing primarily on identifying correct and incorrect answers in LLMs, our study goes further by examining the performance of SLMs across various topics, Elo ratings, and providing a detailed qualitative analysis of the errors. Future work may explore higher-ELO problems, refined prompting strategies, alternative configurations, language-specific generation, and lightweight tuning techniques to further assess and improve model effectiveness.